OpenStudy (ramennoodle01):

what is the identity for sec^2 x - 1/ sin^2 x ?

OpenStudy (anonymous):

\[\frac{ \sec ^2x-1 }{ \sin ^2x }=\frac{ \tan ^2x }{ \sin ^2x }=\frac{ \sin ^2x }{ \cos ^2x }*\frac{ 1 }{ \sin ^2x }=?\]

OpenStudy (ramennoodle01):

thank you so much! I I got the first step right but i couldn't figure out the rest

OpenStudy (anonymous):

yw